You put a flower bed that is 20 by 30ft. you are putting in a uniform border around the outside of the rectangular garden. You have enough to cover 336 sq. feet. How wide should the border be?
Answer by ankor@dixie-net.com(22740)

You put a flower bed that is 20 by 30ft. you are putting in a uniform border
around the outside of the rectangular garden.
You have enough to cover 336 sq. feet.
How wide should the border be?
:
Let x = the width of the borders
:
Then the overall dimensions (garden & border) will be (2x+20) by (2x+30)
:
Overall area - garden area = border area
(2x+20)*(2x+30) - (20*30) = 336
FOIL
4x^2 + 60x + 40x + 600 - 600 = 336
A quadratic equation
4x^2 + 100x - 336 = 0
Simplify, divide by 4
x^2 + 25x - 84 = 0
factors to
(x-3)(x+28) = 0
the positive solution
x = 3 ft is the width of the border
:
:
Check this by finding the overall area (Overall dimensions 26 by 36)
26*36 - 600 =
936 - 600 = 336